OpenStudy (anonymous):

Which pattern follows the given rule? Rule: Start with 2 and then multiply each term by 4 to get the next term. A. 2, 6, 24, 96, . . . B. 2, 6, 10, 14, . . . C. 2, 8, 16, 32, . . . D. 2, 8, 32, 128, . . .

OpenStudy (anonymous):

start = 2 2*4 = a a(4) = b b(4) = c

OpenStudy (anonymous):

is c the anser

OpenStudy (anonymous):

lets test 2*4 = 8 8*4 does not = 8 so no

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

the answer should be D.

OpenStudy (anonymous):

can uhelp me with 2 more please

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

What is the next term in the pattern? Rule: Multiply by 3 and then subtract 1 from the answer to get the next term. 4, 11, , ...

OpenStudy (anonymous):

so they already showed 4 x 3 = 12 - 1 = 11

OpenStudy (anonymous):

right

OpenStudy (anonymous):

Then do 11 x 3 = 33........33 - 1 = 32

OpenStudy (anonymous):

4, 11, 32.......

OpenStudy (anonymous):

You need more?

OpenStudy (anonymous):

so the next number is 42

OpenStudy (anonymous):

no you have to do the process

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so what process do i have oto do 2 get the answer

OpenStudy (anonymous):

32 x 3 = 96...........96 - 1 = 95

OpenStudy (anonymous):

ok then what ?

OpenStudy (anonymous):

ok, i got the answer 95 for the last one. so whatever answer i get you multiply it by 3 and then you subtract 1 from what you got EX: 95 x 3 = 285. then i do, 285 - 1 = 284

OpenStudy (anonymous):

So we got. 4, 11, 32, 95, 284

OpenStudy (anonymous):

Do you have answer choices?

OpenStudy (anonymous):

thats not a choice

OpenStudy (anonymous):

A. 2, 6, 24, 96, . . . B. 2, 6, 10, 14, . . . C. 2, 8, 16, 32, . . . D. 2, 8, 32, 128, . . .

OpenStudy (anonymous):

u there

OpenStudy (anonymous):

ya, sorry

OpenStudy (anonymous):

isnt that the first one you sent

OpenStudy (anonymous):

i was helping you on ur other question you sent